Paul Swanson has an opportunity to acquire a franchise from The Yogurt Place, Inc., to dispense frozen yogurt products underReq 2A Req 2B Req 3A Req 3B Req 1 Prepare a contribution format income statement that shows the expected net operating incomeComplete this question by entering your answers in the tabs below. Req 1 Req 2A Req 2B Req 3A Req 3B Compute the simple rateIf Mr. Swanson requires a simple rate of return of at least 21%, should he acquire the franchise? Yes ONo Req 2B Req 2A Req 3Compute the payback period on the outlet. (Round your answer to 1 decimal place.) years Payback period Req 2B Req 3BIf Mr. Swanson wants a payback of three years or less, will he acquire the franchise? OYes ONo Req 3A Req 3B
Answer #1

1.

Particulars Amount ($) Amount ($)
Sales $   450,000.00
Variable expenses:
Cost of ingredients (450,000*20%) $      90,000.00
Commissions (450,000*15%) $      67,500.00 $   157,500.00
Contribution margin $   292,500.00
Fixed Expenses :  
Salaries $      85,000.00
Rent (4,200*12) $      50,400.00
Depreciation [(360000-24000)/15 years] $      22,400.00
Insurance $        5,000.00
Utilities $      42,000.00 $   204,800.00
Net income $     87,700.00

2A

Simple rate of Return   24.4%
Simple rate of return = Annual incremental net operating income / Initial investment
=87700/360000=24.4%

2B

2B. Yes, because a simple rate of return is more than 21%.

3A

Payback Period 3.3 years
Payback period = Investment required / Annual net cash inflow
= 360,000 / (87700+22400) = 3.3 years
Annual net cash inflow = Net income + Depreciation
3B. It should not acquired as payback period is more than 3 years.